# Ajay Muthukkumar6, Arvind Rajan6, Jaya Kolluri2, Amelia Fiske7

### v1

@@ 1 Introduction
Large language models (LLMs) have forever changed the landscape of natural language processing (NLP), enabling transformative progress in many applications [1, 2, 3, 4, 5]. As LLMs are increasingly applied in diverse domains, the importance of comprehensive benchmarks to assess their performance, limitations, and robustness has become evident. Existing benchmarks such as $GLUE$ [6] and $SuperGLUE$ [7] have been instrumental in evaluating general language understanding and complex task performance. However, the focus of current research challenges has shifted beyond text coherence and linguistic features of language models; instead, challenges of domain knowledge grounding [1, 2, 8, 9, 10, 11], safety [12, 13, 14, 15, 16, 17], hallucinations [18, 19], and bias [20, 21, 22, 23] have emerged as larger issues at hand. These problems are further magnified when extended into high-stakes domains like healthcare, given the potential for biased or inaccurate outputs [24, 25, 26, 27, 28] to influence disparities in health care and outcomes.
This paper investigates the critical issue of representational biases in LLMs, focusing on medical information. Our research explores the interplay between biases in pretraining datasets and their manifestation in language models’ perceptions of disease demographics. We employ quantitative methods to analyze the co-occurrence statistics between demographic subgroups and diseases in $ThePile$ dataset [29].  [^1] Raw counts for subsets of $Arxiv,Github,Wikipedia,StackExchange$ from $RedPajama$ are also available online for subsequent evaluation, but no formal analysis is conducted on these datasets in this study [30]. Subsequently, we assess the correspondence between these associations and the logits produced by various LLMs (across sizes, languages, and alignment methods) for disease-demographic subgroup pairs, offering an in-depth evaluation of how pretraining data biases impact models’ knowledge representation.
Additionally, we benchmark the model-derived associations against actual disease prevalences in the United States among different demographic subgroups. We use this comparison as a grounding mechanism to define the discrepancies between model perceptions and real-world epidemiological data. Our objective is to elucidate the influence of pretraining biases on models’ worldviews, particularly in the critical domain of biomedicine.
@@
Existing bias metrics in the general domain have currently relied on human-annotated examples and focussed on overt stigmatization and prejudices [31, 22, 23]. In contrast, our work examines bias through a different paradigm rooted in real-world data to provide a domain-specific framework for assessing model biases and grounding. We demonstrate this gap using sub-populations defined by United States census categories for gender and race/ethnicity, and normalized disease codes. While these categorizations are necessarily simplistic and imperfect, we contend that the fact that variation is consistently observed across model architectures, model sizes, subgroups, and diseases means that these findings are essential and broadly relevant.
We aim to provide a foundation for future research that evaluates the subgroup robustness of language model associations and equip researchers and practitioners with tools to uncover and understand the biases inherent in their models, thereby facilitating the development of more equitable and effective NLP systems for healthcare. Our full workflow can be found in Figure 1.
Our work makes the following key contributions:
@@
  • We quantitatively analyze the co-occurrences between demographic subgroups and disease keywords in prominent pretraining datasets such as $ThePile$ and release their counts for public use.
@@
  • We examine model logits across various architectures, sizes, and alignment methods using ten template variants that test robustness to disease-demographic subgroup pairs. Our findings illustrate how representational differences in pretraining datasets across diseases align with these logits, regardless of model size and architecture.
  • We benchmark model-derived associations against real-world disease prevalences to highlight discrepancies between model perceptions and actual epidemiological data across model sizes and architectures.
@@
  • We further benchmark these model-derived associations against real-world disease prevalences in different languages (Chinese, English, French, and Spanish) to highlight discrepancies between model perceptions across different languages regardless of model size and alignment methods.
@@
  • Finally, we generate a publicly available web app, www.crosscare.net, that enables easy exploration of these data and the ability to download specific counts, logits, and associations for further work in interpretability, robustness, fairness, and beyond.
@@ 3.1.1 Datasets
3.1.2 Co-occurrence pipeline updates
@@
Future initiatives aim to develop more detailed, consistent, and comprehensive data collection and analysis strategies in collaboration with other research centers.
3.1.3 Validation of Keyword Frequency and Document Co-Occurrence
@@ 3.2.1 Definitions and Variables
3.2.2 Logits Definition
In the context of language models, logits: $z$ refer to the raw output scores from the final layer of the model before any normalization or activation function (such as softmax) is applied. These scores are used to represent the model’s unnormalized prediction probabilities. Given a particular input, logits reflect the model’s preference for each potential output, translating into the predicted probabilities for each class/token set after applying the softmax function.
3.2.3 Logits Calculation
For each model $m$, language $l$, disease $d$, and demographic subgroups $s$, calculate the average logits as follows:
@@ 3.2.3 Logits Calculation
3.2.4 Model’s Disease Demographic Ranking
@@ 3.3 Comparing Rank Order Lists
To understand the representation of diseases across demographic subgroups in different data contexts, we utilized Kendall’s $\tau$ correlation coefficient, which can be defined as:
We computed the Kendall’s tau scores to compare the rank order of diseases based on: Model Logit Ranks versus Pile Demographic Subgroup Co-occurrence Ranks: This comparison assessed how model predictions align with the observed co-occurrence frequencies of disease co-occurrence counts within specific demographic subgroups in the training data. Model Logit Ranks versus Real-World Gold Subset Ranks: This analysis examined the alignment of model outputs with disease rankings derived from a curated real-world dataset, providing insights into the model’s ability to mirror actual disease prevalence across different demographic subgroups.
3.3.1 Variance/Drift in Disease Ranking
@@
3.3.2 Ranking Variance Analysis
@@
Ranking Variance Analysis: First, we calculated Kendall’s tau for each disease across demographic subgroups as previously but instead compared ranks of the base model $m_{0}$ to the ranks of a different, aligned model, $m$. The comparison formula for Kendall’s $\tau$ between the base model and each aligned model is
@@
Here, $s_{i}$ and $s_{j}$ are distinct subgroups, with $i<j$ denoting that each pair of elements is only compared once.
Overall Variance Calculation: Secondly, we computed the average Kendall’s tau for all diseases and demographic subgroups between the two models, evaluating the overall drift from the base model’s ranking:
@@ 3.4 Definition of Controlled and in the Wild
“The controlled group" includes Mamba and Pythia models, which are strictly pre-trained on $ThePile$ only. Here, we aimed to compare these models’ representation of disease prevalence against the real-world prevalence and Pile co-occurrence prevalence.
Additionally, we expanded our evaluation to include “models in the wild," which are publicly accessible and varied in their training and tuning datasets. This group includes base models, such as Llama2, Llama3, Mistral, and Qwen1.5 from the 7b and 70b model sets, and those that have undergone specific alignment methods, including RLHF [67], SFT [68], or DPO [69], and also biomedical domain-specific continued pre-training (detailed models’ descriptions at Appendix C Table 4). We accessed their model logits with four languages (English, Spanish, French, Chinese). This dual approach of controlled evaluation and real-world model assessment allowed for a comprehensive analysis of models’ understanding of disease’s real-world prevalence across languages and how alignment methods might alter it.
3.4.1 Experimental Framework
@@ 3.5 Findings
3.5.1 Variation Across Windows
We evaluated the ranks across different token window sizes of 50, 100, and 250 within each disease demographic pair. No difference was observed in the top disease rank across each window size’s ranking. For the remainder of the paper, we use the 250-token window for simplicity, but the raw counts across each window size for each disease are available at https://crosscare.net/downloads.
3.5.2 Demographic Distributions
We collected all 89 disease co-occurrences in $ThePile$ and 15 real-world prevalences from CDC (Appendix A Table 2). Within both $ThePile$ datasets, White was the most frequently represented race/ethnicity subgroup (87/89), most commonly followed by Black and Hispanic subgroups with relatively lower counts. The least represented race/ethnicity subgroups were consistently Pacific Islanders and Indigenous. However, among real-world statistics, Indigenous is often the top-ranked subgroup, followed by white and Black subgroups.
@@ 3.5.2 Demographic Distributions
For gender distribution in $ThePile$, the male subgroup was more prevalent than the female subgroup for the reported diseases, with the non-binary subgroup being the least represented (Appendix A Table 2).
Figure 2 shows demographic subgroup rank according to real-world prevalence, $ThePile$ co-occurrence counts, and Llama3 logits for the 15 diseases where real-world prevalence is available. This shows discrepancies and alignments between dataset co-occurrence representations and actual demographic prevalence of diseases. The raw counts and ranking in $ThePile$ dataset versus real-world prevalence from the NHIS survey are further elaborated in Appendix A Table 3.
@@ 4.1 Models in the Controlled Group
4.1.1 Logits Rank vs Co-occurrence
@@
Figure 3: Top ranked gender and race/ethnicity subgroups across each of the 89 diseases and the suite of Pythia and Mamba models according to logits results (stacked bars). Co-occurrence and logit rank match demonstrate the number of diseases for which the top-ranked demographic subgroup is the same when calculated using co-occurrences and logits (black line). Demographic subgroups that did not appear as the top-ranked group are not shown.
4.1.2 Logits Rank vs Co-occurrence vs Real Prevalence
@@
Figure 4: Kendall’s tau of Mamba and Pythia’s logits vs co-occurrence, and real prevalence for gender (top) and race/ethnicity (bottom).
4.1.3 Rank vs Co-occurrence counts
@@ 4.2 Models in the wild
For all models that we tested across size, alignment method, and language, no model’s disease logits rankings had $\tau>0.35$ (Min = -0.6, Max = 0.32, Median = -0.07, Avg = -0.06, Var = 0.03) for gender or race/ethnicity, suggesting none had good knowledge of real-world prevalence. Figure 2 illustrates discrepancies between Llama3’s logits compared to $ThePile$ and real-world prevalences. These discrepancies might lead to incorrect and/or biased judgments in healthcare settings.
4.2.1 Variation across Alignment strategies
@@
The impact of different alignment strategies on the LLama2 70b series for both race/ethnicity and gender are displayed in Figure 5. None of the alignment methods nor in-domain continued pre-training corrected the base model towards more accurate reflections of real-world prevalence. In fact, we observed some of the debiasing strategies during alignment adversely impacting the model’s decisions (Table 1). All Llama2 70b series alignment methods increased preference for female over male subgroups, and decreased preference for the Black subgroup, in English. A similar observation was seen among the Mistral family. For Qwen1.5-7b base compared to Qwen1.5-7b chat in English, PPO+DPO shifted its favor to the Indigenous instead of Asian subgroup (Appendix C.2.3 Table 6).
@@
For the Llama2 70b series, models tuned by different alignment methods (SFT, DPO) did not change the rank-ordering of race/ethnicity subgroups ($\delta>=0.8$). Models that demonstrated noticeable variation were the Meditron variant, which underwent continued pre-training on medical domain data, and the chat version that went through reinforcement learning with human feedback (RLHF). Similar trends were observed for Mistral’s gender results, where Bio-mistral was given continued pre-training on biomedical text with Mistral-instruct with RLHF (Appendix C.2.1 Table 5(b)).
Language Model Name Alignment A B H I PI W $\delta\uparrow$ $\tau\uparrow$
English Llama2-70b Base 0 70 0 0 0 19 N/A -0.11
Llama2-70b chat RLHF 0 86 0 0 0 3 0.47 -0.21
Tulu2-70b SFT 0 65 0 0 0 24 0.87 -0.13
Tulu2-70b-dpo DPO 1 54 0 0 0 34 0.87 -0.13
Meditron-70b Biomed 0 60 0 0 0 29 0.65 -0.19
Chinese Llama2-70b Base 0 0 0 0 0 89 N/A -0.13
Llama2-70b chat RLHF 0 0 0 0 0 89 0.93 -0.07
Tulu2-70b SFT 0 0 0 0 0 89 0.99 -0.15
Tulu2-70b-dpo DPO 0 0 0 0 0 89 0.97 -0.16
Meditron-70b Biomed 0 0 0 0 0 89 1.0 -0.13
Spanish Llama2-70b Base 0 89 0 0 0 0 N/A 0.04
Llama2-70b chat RLHF 0 89 0 0 0 0 1.0 0.04
Tulu2-70b SFT 0 89 0 0 0 0 1.0 0.04
Tulu2-70b-dpo DPO 0 89 0 0 0 0 1.0 0.04
Meditron-70b Biomed 0 89 0 0 0 0 0.91 0.08
French Llama2-70b Base 0 0 0 0 0 89 N/A -0.16
Llama2-70b chat RLHF 0 0 0 0 0 89 0.02 0.07
Tulu2-70b SFT 0 0 0 0 0 89 0.88 -0.15
Tulu2-70b-dpo DPO 0 0 0 0 0 89 0.88 -0.15
Meditron-70b Biomed 0 0 0 0 0 89 0.48 -0.23
(a) Demographic distribution by race and model alignment
English Llama2-70b Base 36 53 0 N/A -0.47
Llama2-70b chat RLHF 19 70 0 0.77 -0.07
Tulu2-70b SFT 22 67 0 0.85 -0.20
Tulu2-70b-dpo DPO 10 79 0 0.79 -0.07
Meditron-70b Biomed 16 73 0 0.79 -0.07
Chinese Llama2-70b Base 88 1 0 N/A -0.07
Llama2-70b chat RLHF 1 0 88 -0.12 -0.47
Tulu2-70b SFT 53 34 2 0.61 -0.60
Tulu2-70b-dpo DPO 49 39 1 0.6 -0.60
Meditron-70b Biomed 75 14 0 0.9 -0.33
Spanish Llama2-70b Base 44 0 45 N/A 0.07
Llama2-70b chat RLHF 83 0 6 0.68 0.07
Tulu2-70b SFT 85 0 4 0.69 0.07
Tulu2-70b-dpo DPO 87 0 2 0.68 0.07
Meditron-70b Biomed 89 0 0 0.66 0.07
French Llama2-70b Base 0 89 0 N/A -0.07
Llama2-70b chat RLHF 9 80 0 0.93 -0.33
Tulu2-70b SFT 0 89 0 1.0 -0.07
Tulu2-70b-dpo DPO 0 89 0 1.0 -0.07
Meditron-70b Biomed 0 89 0 1.0 -0.07
(b) Demographic distribution by gender and model alignment
$\tau:\{-1,1\}$ Kendal Tau of model’s prevalence representation vs real-world prevalence
4.2.2 Models’ representation across different languages
@@
We also observed differences in models’ representation across languages as shown in Figure 5. For all Mistral and Llama series models, we observed a preference toward female subgroup in Chinese but male subgroup in French (Appendix C.2.1 Table 5(b)). In Qwen, we observed an overall preference towards male subgroup in Chinese, Spanish, and French but female subgroup in English (Appendix C.2.3 Table 6).
@@
For race/ethnicity, templates in English and Spanish showed preference for the Black subgroup, and templates in Chinese and French showed preference for the white subgroup (Llama2 series at Table 1, Mistral at Appendix C.2.1 Table 5(b)). Interestingly, for the series of Qwen1.5 models, which were pretrained on mostly English and Chinese, we observed a strong bias toward Asian subgroup in Chinese and English templates, and Black subgroup in Spanish and French templates (Appendix C.2.3 Table 6).

### v2

@@ 1 Introduction
Large language models (LLMs) enabled transformative progress in many applications [1, 2, 3, 4, 5]. Benchmarks to assess language models, such as $GLUE$ [6] and $SuperGLUE$ [7], are instrumental in evaluating general language understanding and complex task performance. However, as LLMs are increasingly applied in diverse domains, challenges of domain knowledge grounding [1, 2, 8, 9, 10, 11], safety [12, 13, 14, 15, 16, 17], hallucinations [18, 19], and bias [20, 21, 22, 23] have emerged as important issues that are inadequately assessed by existing benchmarks. These problems magnified in high-stakes domains like healthcare, given the potential for biased or inaccurate outputs [24, 25, 26, 27, 28] to influence disparities in health care and outcomes.
@@
This paper investigates representational biases in LLMs, focusing on medical information. Our research explores the interplay between biases in pretraining datasets and their manifestation in LLMs’ perceptions of disease demographics. Existing bias metrics in the general domain have currently relied on human-annotated examples and focused on overt stigmatization and prejudices [31, 22, 23]. In contrast, our work examines bias through a different paradigm rooted in real-world data to provide a domain-specific framework for assessing model biases and grounding. We demonstrate this gap using sub-populations defined by United States census categories for gender and race/ethnicity, and normalized disease codes. While these categorizations are necessarily simplistic and imperfect, we contend that the fact that variation is consistently observed across model architectures, model sizes, subgroups, and diseases means that these findings are meaningful and broadly relevant. We aim to provide a foundation for future research that evaluates the subgroup robustness of LLM associations and equip researchers and practitioners with tools to uncover and understand the biases inherent in their models, thereby facilitating the development of more equitable and effective NLP systems for healthcare. Our full workflow can be found in Figure 1.
Specifically, our work makes the following key contributions:
@@
1. We conduct a quantitative analysis of the co-occurrences between demographic subgroups and disease keywords in prominent pretraining datasets like The Pile, releasing their counts publicly.
@@
2. We evaluate model logits across various architectures, sizes, and alignment methods using ten template variants to test robustness to disease-demographic subgroup pairs. Our findings reveal that representational differences in pretraining datasets across diseases align with these logits, irrespective of model size and architecture.
@@
3. We benchmark model-derived associations against real-world disease prevalences to highlight discrepancies between model perceptions and actual epidemiological data. Additionally, we compare these associations across different languages (Chinese, English, French, and Spanish) to emphasize discrepancies across languages.
@@
4. We provide a publicly accessible web app, www.crosscare.net, for exploring these data and downloading specific counts, logits, and associations for further research in interpretability, robustness, and fairness.
@@ 3.1 Methods
Co-occurrence pipeline updates
@@
Validation of Keyword Frequency and Document Co-Occurrence
@@ 3.2 Mathematical Description of Prevalence Calculation Using Average Logits
Logits Definition
In the context of language models, logits: $z$ refer to the raw output scores from the final layer of the model before any normalization or activation function (such as softmax) is applied. These scores are used to represent the model’s unnormalized prediction probabilities. Given a particular input, logits reflect the model’s preference for each potential output, translating into the predicted probabilities for each class/token set after applying the softmax function. For each model $m$, language $l$, disease $d$, and demographic subgroups $s$, calculate the average logits as follows:
@@
Model’s Disease Demographic Ranking
@@ 3.3 Comparing Rank Order Lists
We utilized Kendall’s $\tau$ correlation coefficient to understand the representation of diseases across demographic subgroups in different data contexts here (see details in Appendix A.2).
Variance/Drift in Disease Ranking
@@
Ranking Variance Analysis
@@
First, we calculated Kendall’s tau for each disease across demographic subgroups as previously but instead compared ranks of the base model $m_{0}$ to the ranks of a different, aligned model, $m$. The comparison formula for Kendall’s $\tau$ between the base model and each aligned model is
@@
Here, $s_{i}$ and $s_{j}$ are distinct subgroups, with $i<j$ denoting that each pair of elements is only compared once. Secondly, we computed the average Kendall’s tau for all diseases and demographic subgroups between the two models, evaluating the overall drift from the base model’s ranking:
@@ 3.4 Definition of Controlled and in the Wild
“The controlled group” includes Mamba and Pythia models, which are strictly pre-trained on $ThePile$ only. Here, we aimed to compare these models’ representation of disease prevalence against the real-world prevalence and Pile co-occurrence prevalence.
Additionally, we expanded our evaluation to include “models in the wild”, which are publicly accessible and varied in their training and tuning datasets. This group includes base models, such as Llama2, Llama3, Mistral, and Qwen1.5 from the 7b and 70b model sets, and those that have undergone specific alignment methods, including RLHF [67], SFT [68], or DPO [69], and also biomedical domain-specific continued pre-training (detailed models’ descriptions at Appendix C.1 Table 3). We accessed their model logits with four languages (English, Spanish, French, Chinese). This dual approach of controlled evaluation and real-world model assessment allowed for a comprehensive analysis of models’ understanding of disease’s real-world prevalence across languages and how alignment methods might alter it.
Experimental Framework
@@ 3.5 Findings
Variation Across Windows
We evaluated the ranks across different token window sizes of 50, 100, and 250 within each disease demographic pair. No difference was observed in the top disease rank across each window size’s ranking. For the remainder of the paper, we use the 250-token window for simplicity, but the raw counts across each window size for each disease are available on our website.
Demographic Distributions
We collected all 89 disease co-occurrences in $ThePile$ and 15 real-world prevalences from CDC (Appendix A.3 Table 1). Within both $ThePile$ datasets, White was the most frequently represented race/ethnicity subgroup (87/89), most commonly followed by Black and Hispanic subgroups with relatively lower counts. The least represented race/ethnicity subgroups were consistently Pacific Islanders and Indigenous. However, among real-world statistics, Indigenous is often the top-ranked subgroup, followed by white and Black subgroups.
@@
For gender distribution in $ThePile$, the male subgroup was more prevalent than the female subgroup for the reported diseases, with the non-binary subgroup being the least represented (Appendix A.3 Table 1).
Figure 2 shows demographic subgroup rank according to real-world prevalence, $ThePile$ co-occurrence counts, and Llama3 logits for the 15 diseases where real-world prevalence is available. This shows discrepancies and alignments between dataset co-occurrence representations and actual demographic prevalence of diseases. The raw counts and ranking in $ThePile$ dataset versus real-world prevalence from the NHIS survey are further elaborated in Appendix A.3 Table 2.
@@ 4.1 Models in the Controlled Group
Logits Rank vs Co-occurrence
@@
Figure 3: a) Top ranked gender (top) and race/ethnicity (bottom) subgroups across 89 diseases and the suite of Pythia and Mamba models according to logits results (stacked bars). Co-occurrence and logit rank match demonstrate the number of diseases for which the top-ranked demographic subgroup is the same when calculated using co-occurrences and logits (black line). Demographic subgroups that did not appear as the top-ranked group are not shown. b) Kendall’s tau of Mamba and Pythia’s logits vs co-occurrence, and real prevalence for gender (top) and race/ethnicity (bottom).
Logits Rank vs Co-occurrence vs Real Prevalence
@@
Rank vs Co-occurrence counts
@@ 4.2 Models in the wild
For all models that we tested across size, alignment method, and language, no model’s disease logits rankings had $\tau>0.35$ (Min = -0.73, Max = 0.33, Median = -0.05, Avg = -0.06, Var = 0.03) for gender or race/ethnicity, suggesting none had good knowledge of real-world prevalence. Figure 2 illustrates discrepancies between Llama3’s logits compared to $ThePile$ and real-world prevalences. These discrepancies might lead to incorrect and/or biased judgments in healthcare settings.
Variation across Alignment strategies
@@
The impact of different alignment strategies on the LLama2 70b series for both race/ethnicity and gender are displayed in Figure 4. None of the alignment methods nor in-domain continued pre-training corrected the base model towards more accurate reflections of real-world prevalence. In fact, we observed some of the debiasing strategies during alignment adversely impacting the model’s decisions (Appendix C.2.3 Table 6(b)). All Llama2 70b series alignment methods increased preference for female over male subgroups, and decreased preference for the Black subgroup, in English. A similar observation was seen among the Mistral family. For Qwen1.5-7b base compared to Qwen1.5-7b chat in English, PPO+DPO shifted its favor to the Indigenous instead of Asian subgroup (Appendix C.2.2 Table 5(b)).
@@
For the Llama2 70b series, models tuned by different alignment methods (SFT, DPO) did not change the rank-ordering of race/ethnicity subgroups ($\delta>=0.8$). Models that demonstrated noticeable variation were the Meditron variant, which underwent continued pre-training on medical domain data, and the chat version that went through reinforcement learning with human feedback (RLHF). Similar trends were observed for Mistral’s gender results, where Bio-mistral was given continued pre-training on biomedical text with Mistral-instruct with RLHF (Appendix C.2.1 Table 4).
Figure 4: Top ranked gender and race/ethnicity subgroups across each of the 89 diseasese and different alignments of methods for Llama2 models according to logits results (stacked bars).
Models’ representation across different languages
@@
We also observed differences in models’ representation across languages as shown in Figure 4. For all Mistral and Llama series models, we observed a preference toward female subgroup in Chinese but male subgroup in French (Appendix C.2.1 Table 4, Appendix C.2.3 Table 6(b)). In Qwen, we observed an overall preference towards male subgroup in Chinese, Spanish, and French but female subgroup in English (Appendix C.2.2 Table 5(b)).
@@
For race/ethnicity, templates in English and Spanish showed preference for the Black subgroup, and templates in Chinese and French showed preference for the white subgroup (Llama2 series at Appendix C.2.3 Table 6(b), Mistral at Appendix C.2.1 Table 4). Interestingly, for the series of Qwen1.5 models, which were pretrained on mostly English and Chinese, we observed a strong bias toward Asian subgroup in Chinese and English templates, and Black subgroup in Spanish and French templates (Appendix C.2.2 Table 5(b)).